OpenStudy (anonymous):

Consider the experiment of tossing a loaded coin in which heads is three times as likely as tails to occur. Simulation of the experiment uses a random number function, considering a toss to be tails (T) if the result is less than 0.25, and considering a toss to be heads (H) if the result is greater than or equal to 0.25. The result of 10 simulations is shown below. H, H, H, T, H, H, H, T, H, T Using these 10 outcomes of the experiment estimate the probabilities P(H) and P(T) by the ratios. P(H) = number of times H occurred/10 =??? P(T) = number of times T occurred/10 =??

OpenStudy (anonymous):

what are the actual probabilities? P(H) =? P(T) =?

OpenStudy (anonymous):

P(h)=.7 P(T)-.3

OpenStudy (anonymous):

*P(T)=.3

OpenStudy (anonymous):

what about the number of times tales occurred divided by 10?

OpenStudy (anonymous):

obviously its the same number but they said it was incorrect last time i put that

OpenStudy (anonymous):

anyone know?

OpenStudy (dumbcow):

P(H) = 7/10 P(T) = 3/10

OpenStudy (anonymous):

but that wasn't the correct answer.....

OpenStudy (dumbcow):

is there any other part of the question? They want P(H) and P(T) based on the simulation correct?

OpenStudy (anonymous):

yes correct which is why i do not understand why that is not the correct answer.....

OpenStudy (dumbcow):

hmm me neither

OpenStudy (anonymous):

Although the sample shows differently, techincally, the proability of heads is 1/2 and tails is 1/2 ... Do those input as the correct answer?

OpenStudy (anonymous):

they said it has something to do with theheads being three times more likely to occur

OpenStudy (anonymous):

the actual probabilities were correct just not the estimated ones

OpenStudy (dumbcow):

Its a loaded coin...where the actual probabilities are P(H) = 0.75 P(T) = 0.25

OpenStudy (anonymous):

wait i lied. the actual probabilities were incorrect

OpenStudy (anonymous):

Ahh, see you solved it yourself.. I guess it would've helped if I read the problem.. ;/

OpenStudy (anonymous):

nvm...?

OpenStudy (dumbcow):

The estimated probabilities must be P(H) = 7/10 P(T) = 3/10

OpenStudy (anonymous):

yes that is correct. now what are the actual probabilities?

OpenStudy (dumbcow):

see above ^^

OpenStudy (anonymous):

If it is 3 times more likely, Acutual probabilities are: p(h) = 3/4 p(t) = 1/4